# An Aligning and Training Framework for Multimodal Recommendations

### v1

@@ Abstract.
With the development of multimedia applications, multimodal recommendations are playing an essential role, as they can leverage rich contexts beyond user interactions. Existing methods mainly regard multimodal information as an auxiliary, using them to help learn ID features; however, there exist semantic gaps among multimodal content features and ID features, for which directly using multimodal information as an auxiliary would lead to misalignment in representations of users and items. In this paper, we first systematically investigate the misalignment issue in multimodal recommendations, and propose a solution named AlignRec. In AlignRec, the recommendation objective is decomposed into three alignments, namely alignment within contents, alignment between content and categorical ID, and alignment between users and items. Each alignment is characterized by a specific objective function and is integrated into our multimodal recommendation framework. To effectively train our AlignRec, we propose starting from pre-training the first alignment to obtain unified multimodal features and subsequently training the following two alignments together with these features as input. As it is essential to analyze whether each multimodal feature helps in training, we design three new classes of metrics to evaluate intermediate performance. Our extensive experiments on three real-world datasets consistently verify the superiority of AlignRec compared to nine baselines. We also find that the multimodal features generated by AlignRec are better than currently used ones, which are to be open-sourced.
*Yifan Liu and Kangning Zhang contributed equally to this research. The email for Yifan Liu is ¡sjtulyf123@sjtu.edu.cn¿.
copyright: acmcopyrightjournalyear: 2024doi: 10.1145/1122445.1122456conference: KDD ’24: In Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining; August 25–29, 2024; Barcelona, Spainbooktitle: KDD ’24: In Proceedings of the 30th ACM SIGKDD Conference on Knowledge Discovery and Data Mining, August 25–29, 2024, Barcelona, Spainprice: 15.00isbn: 978-1-4503-XXXX-X/18/06
@@ 1. Introduction
Recommender systems have become one of the essential sources for users to explore content from online services such as e-commerce platforms and social media, especially when ID features are helpless in long-tail items or cold-start scenarios. Several works on multimodal recommendation have demonstrated that utilizing multimodal information (e.g., images and texts) together with categorical features (e.g., itemID and category) can gain great success in multiple recommendation challenges (He and McAuley, 2016b; Wei et al., 2019, 2020; Zhou et al., 2023; Zhou and Shen, 2023; Yu et al., 2023).
Meanwhile, alongside the development of recommender systems, the field has witnessed significant advancements in multimodal large language models (mLLMs) like GPT-4V (Achiam et al., 2023) and Gemini (Team et al., 2023). One of the key lessons towards generalized mLLMs is the alignment among different content modalities, i.e., viewing different content modalities residing in a unified space (Li et al., 2021; Lin et al., 2023; Girdhar et al., 2023; Wang et al., 2022), because the distribution gaps between different modalities’ representations will significantly harm the cross-modality understanding and make it hard to extract and utilize multimodal features. However, the alignment issue has not been well-studied in the area of multimodal recommendation. Early works (He and McAuley, 2016b; Wei et al., 2019, 2020) view multimodal information as an auxiliary feature and concatenate it to categorical features, ignoring the distribution gap between multimodal features and categorical features, since they are modeled for different objects. Most recently, researchers explore auxiliary losses with multimodal information for deriving better categorical embeddings via self-supervised learning (Tao et al., 2022; Zhou et al., 2023) or graph structures (Zhou and Shen, 2023; Yu et al., 2023), which still suffer from the misalignment across content modalities.
Figure 1. Motivations of alignment. Left is the t-SNE (Van der Maaten and Hinton, 2008) results of vision and text features released in Amazon Sports dataset. Right is the t-SNE results of well-trained content and ID modality feature pairs using VBPR (He and McAuley, 2016b) method.
@@
Challenge 1: How to perform alignment in multimodal recommendation scenario has not been carefully designed. The input of multimodal recommendation contains not only content modalities like images and text, but also various ID features; And they all remain on their own spaces with different distributions. We illustrate the t-SNE results of some features in Amazon Sports dataset in Fig. 1, discovering that the image and text modality have quite different distributions, and the content and ID feature pair of the same item can also be far away [^1] More analysis on current methods and cases are in Appendix A.. Therefore, a unified design for aligning all modalities is required to handle misalignment.
@@
Challenge 2: The popular end-to-end training with joint optimization on alignment and recommendation tasks is hard to learn well. The learning speed of learning content modalities and the categorical features are not consistent. Specifically, the former working on images and texts often requires a large amount of data and time to learn sufficient semantic knowledge (Li et al., 2021), whereas the latter only needs a few epochs (Zhang et al., 2022). Therefore, the speed of learning content modalities is significantly slower than learning categorical modalities. Joint training for multimodal recommendation is hard to deal with this imbalanced rate, introduces high computational overhead, and finally leads to an insufficient representation of items.
@@
Challenge 3: The influence of content modality to multimodal recommendation has not been fully studied. There are a lot of metrics designed for mLLMs to determine whether the aligned features are suitable for downstream tasks (Radford et al., 2021). However, the multimodal recommendation just incorporates the vision and text features into recommendation and evaluates the final performance, which is indirect and does not consider the quality of these features. If the vision or text features are not suitable for recommendation, more effort is required for re-generation, decreasing the recommendation efficiency. We utilize a zero-shot evaluation protocol (see detail in Sec. 4.4.2) on Amazon Baby dataset, and surprisingly find that the absence of vision features can gain a relatively 12.1% increase of recall@20. These observations call for us to directly evaluate the influence of multimodal features to recommendation and produce more suitable features from the data science perspective.
@@
To manage the challenges above, we propose a multimodal recommendation framework AlignRec. For Challenge 1, we develop three alignment objectives for different contexts, respectively. The alignment within contents, namely inter-content alignment (ICA), aligns different content modality fields (e.g., vision and text) with the help of attention-based cross-modality encoder, and outputs a unified modality representation for one item. The alignment between content and categorical ID, namely content-category alignment (CCA), utilizes contrastive learning to bridge the gap between multimodal content features and user/item ID-based features for better utilization and adaptation of multimodal information. The user-item alignment (UIA) aligns users and their interacted items via cosine similarity. For Challenge 2, we propose to train AlignRec by first pre-training the inter-content alignment task, and then training the remaining two alignment tasks together with the recommendation goal. With the pre-training stage utilizing mLLM’s capacity to align content modality, the process of jointly learning multi-modality becomes more stable, which leads to a better item/user representation. For Challenge 3, we propose three intermediate evaluation protocols, including zero-shot, item-CF and mask modality recommendation, to directly evaluate whether the unified and aligned multimodal features generated in the pre-training process via inter-content alignment are helpful for recommendation. Our contributions can be concluded as follows:
@@
  • We also provide three classes of new intermediate protocols to evaluate whether the multimodal features are effective or not. To the best of our knowledge, we are the first to intermediately evaluate the multimodal recommendation.
We conduct experiments on three public datasets and show the state-of-the-art performance of AlignRec compared to nine baselines. The intermediate evaluation with in-depth analysis also shows the superior performance of our alignments and training strategies. Our proceeded multimodal features showing better performance than existing ones will also be released upon publication.
@@ 2. Related Work
Multimodal Recommendation. Multimodal recommendation is derived from traditional collaborative filtering (CF) based recommendation paradigm (Su and Khoshgoftaar, 2009), and incorporates multimodal information for better recommendation. It is an extension from matrix factorization methods (Koren et al., 2009) and CF methods like BPR (Rendle et al., 2012) and LightGCN (He et al., 2020a) and is usually effective in cold-start or long-tail scenarios.
@@
Early multimodal recommendation methods only concatenate vision or text features with ID-based features. VBPR (He and McAuley, 2016b) extends BPR-based recommendation by adding visual dimensions to neural networks, and its visual features are generated from pre-trained CNNs. ACF (Chen et al., 2017) applies an attention layer to calculate modalities’ contributions to items. MMGCN (Wei et al., 2019) and GRCN (Wei et al., 2020) build modality graphs and use GCN (Kipf and Welling, 2016) to aggregate their information before concatenated with ID embeddings. These methods ignore the misalignment between ID-based features and content features, and thus cannot fully exploit multimodal information into recommendation.
@@
Recently, some methods make efforts to shorten the distance between item’s ID and its content information. MMGCL (Yi et al., 2022) applies negative sampling to discover modalities relationship. SLMRec (Tao et al., 2022) builds self-supervised tasks like feature drop to enhance the modality representation. BM3 (Zhou et al., 2023) builds text and vision alignment loss by contrastive view generator. These methods only pay attention to the content-category alignment. Besides, optimizing them together is hard to converge, leading to an insufficient representation.
@@
Aside from utilizing losses for multimodal information, some methods try to fuse vision/text information manually. LATTICE (Zhang et al., 2021) and FREEDOM (Zhou and Shen, 2023) construct modality-similarity graphs using multimodal features and combine these graphs with an importance weight to get fused representations. MGCN (Yu et al., 2023) adopts attention layers to capture the modalities’ importance. These methods do not explicitly model the alignment, and will suffer from the modality incompatibility problem (e.g., conflict between text and vision).
@@
Vision-Language Pre-training. Recently, vision-language pre-training is effective in understanding and integrating vision and language information with large-scale data, which is helpful in multimodal recommendation. CLIP (Radford et al., 2021) proposes to align image and text via contrastive learning and large-scale image-text pair pre-training. ALBEF (Li et al., 2021) designs an image-text contrastive learning task to align image and text features before fusing them. BLIP (Li et al., 2022) utilizes multi-task learning by building the image-text contrastive learning, image-text matching, and mask language modeling tasks. BEiT3 (Wang et al., 2022) regards image as a foreign language and builds a general make-vision-language task to learn one integrated representation for multimodal features. Vision and language information is naturally fused and aligned with the help of multiway transformers (Bao et al., 2022). They inspire us to align vision and text.
Figure 2. An overview of AlignRec. Part A, B, C is the detailed architectures, with user ID, item ID, text and image as input, and user and item representations as output. Part D, E, F is the corresponding alignment tasks specifically designed in AlignRec.
@@ 3. Problem Formulation
where, $u,i$ is a user-item pair satisfying $r_{u,i}=1$, $i^{\prime}$ is sampled from $\mathcal{I}$ satisfying $r_{u,i^{\prime}}=0$. $f:u\rightarrow\mathbb{R}^{d_{h}}$ and $g:i\rightarrow\mathbb{R}^{d_{h}}$ are mapping functions to map raw features (ID, images, text, etc.) into dense representations with dimension $d_{h}$; $s$ is a score function (e.g., inner product); $\sigma$ is the sigmoid function. Intuitively, for a user $u$, a clicked item $i$, and an un-clicked item $i^{\prime}$, BPR tries to maximize the gap between the clicked u-i score and the un-clicked u-i score.
@@ 4.1. Framework Overview
The overview of AlignRec is illustrated in Fig. 2. The whole framework consists of three alignments with an intermediate evaluation module. In this section, we will first introduce the architectures, then describe our three alignment objectives, and finally introduce training strategies and intermediate evaluation protocols.
@@ 4.2. Architecture Design
For generalization, AlignRec only contains three necessary modules: The multimodal encoder module $MMEnc$ aligns the vision and text modality knowledge into one unified representation. The aggregation module $Aggregator$ aggregates the neighbors’ information with the help of $\mathcal{G}$. The fusion module $Fuser$ fuses multimodal and ID-based representations for recommendation. The goal is to generate aligned item’s and user’s representations $\mathbf{h}_{i}$ and $\mathbf{h}_{u}$ for top-K recommendation.
@@ 4.2.1. Multimodal Encoder Module
The multimodal encoder module aims at aligning vision and text of the same item, and outputs a unified multimodal representation $\textbf{h}_{enc}^{i}$. We propose to adopt a transformer-based multimodal encoder to generate the unified multimodal representation by aligning vision and text information. Formally, suppose $MMEnc$ is a multimodal encoder, and $h_{enc}^{i}$ is the unified representation for item $i$’s vision and text, then:
@@
In Eq. (2), the vision-text alignment is optimized through the cross-attention based transformers in $MMEnc$. In implementations, we utilize the most advanced BEiT3 (Wang et al., 2022) as the backbone. The $MMEnc$ will align vision and text information via cross-modality experts (Bao et al., 2022) and transformers. We train this module in Sec. 4.3.1 through the mask-then-predict strategy to ensure that the model has a better understanding of both vision and text modalities.
@@ 4.2.3. Fusion Module
The multimodal hidden representations and ID-based representations contain the information of item/user from different fields or aspects. In our AlignRec framework, we use a fusion module $Fuser$ to fuse this information into general item/user representations $\mathbf{h}^{i}$ and $\mathbf{h}^{u}$ in Eq. (7).
@@
Specifically, we utilize the element-wise addition operation as presented in Eq. (8) to fuse the multimodal and ID-based representation, and finally get a general representation for users and items. This module can also be replaced by more complicated fusion operations like attention in current methods.

### v4

# AlignRec: Aligning and Training in Multimodal Recommendations

@@ Abstract.
With the development of multimedia systems, multimodal recommendations are playing an essential role, as they can leverage rich contexts beyond interactions. Existing methods mainly regard multimodal information as an auxiliary, using them to help learn ID features; However, there exist semantic gaps among multimodal content features and ID-based features, for which directly using multimodal information as an auxiliary would lead to misalignment in representations of users and items. In this paper, we first systematically investigate the misalignment issue in multimodal recommendations, and propose a solution named AlignRec. In AlignRec, the recommendation objective is decomposed into three alignments, namely alignment within contents, alignment between content and categorical ID, and alignment between users and items. Each alignment is characterized by a specific objective function and is integrated into our multimodal recommendation framework. To effectively train AlignRec, we propose starting from pre-training the first alignment to obtain unified multimodal features and subsequently training the following two alignments together with these features as input. As it is essential to analyze whether each multimodal feature helps in training and accelerate the iteration cycle of recommendation models, we design three new classes of metrics to evaluate intermediate performance. Our extensive experiments on three real-world datasets consistently verify the superiority of AlignRec compared to nine baselines. We also find that the multimodal features generated by AlignRec are better than currently used ones, which are to be open-sourced in our repository https://github.com/sjtulyf123/AlignRec_CIKM24.
Multimodal Recommendation; Representation Alignment; Vision Language Pre-training
price: 15.00isbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems
@@ 1. Introduction
Multimodal recommendations have become one of the essential sources for users to explore content from online services such as social media, especially when ID features are helpless in long-tail items or cold-start scenarios. Several works have demonstrated that utilizing multimodal information (e.g., images and texts) together with categorical features (e.g., ID and category) can gain great success in multiple scenarios (He and McAuley, 2016b; Zhang et al., 2024; Wei et al., 2020; Zhou et al., 2023; Zhou and Shen, 2023; Yu et al., 2023).
Meanwhile, alongside the development of recommender systems, the field has witnessed significant advancements in multimodal large language models (mLLMs) like GPT-4V (Achiam et al., 2023) and Gemini (Team et al., 2023). One of the key lessons towards generalized mLLMs is the alignment among different content modalities, i.e., viewing different content modalities residing in a unified space (Li et al., 2021; Lin et al., 2023; Girdhar et al., 2023; Wang et al., 2022), because the distribution gaps among modalities’ representations will significantly harm the cross-modality understanding and make it hard to utilize multimodal features. However, the alignment issue has not been well-studied in multimodal recommendations. Early works like VBPR (He and McAuley, 2016b) in Fig. 1(a) view multimodal information as an auxiliary feature and concatenate it to categorical features (Wei et al., 2019, 2020), ignoring the distribution gap between multimodal features and categorical features. Some methods (Zhou and Shen, 2023; Yu et al., 2023) explore auxiliary losses with multimodal information for deriving better categorical embeddings via graph structures like FREEDOM (Zhou and Shen, 2023) in Fig. 1(b). But they ignore the inconsistency between image and text. Some self-supervised learning methods (Tao et al., 2022; Zhou et al., 2023) like BM3 (Zhou et al., 2023) in Fig. 1(c) train inter-content loss and recommendation-related user-item loss together, leading to an insufficient representation on items’ content
Figure 1. Comparison among VBPR (He and McAuley, 2016b), FREEDOM (Zhou and Shen, 2023), BM3 (Zhou et al., 2023) and AlignRec. $\mathcal{L}_{ICA},\mathcal{L}_{CCA},\mathcal{L}_{UIA}$ are losses for inter-content alignment, content-category alignment and user-item alignment. Dashed lines are scopes of alignment losses.
@@
Challenge 1: How to perform alignment in multimodal recommendation has not been carefully designed. The input of multimodal recommendation contains not only content modalities like images and text, but also various ID features; And they all remain on their own spaces with different distributions. For content modalities like image and text, they have different semantic information with their own distributions. In experiments, we find that the content and ID feature pair of the same item can also be far away in current methods, which will be discussed in Fig. 4. Therefore, a unified design for aligning all modalities is required to handle misalignment.
@@
Challenge 2: The popular end-to-end training with joint optimization on alignment and recommendation tasks is hard to learn well. The speeds of learning content modalities and the categorical features are not consistent. Specifically, the former working on images and texts often requires a large amount of data and time to learn sufficient semantic knowledge (Li et al., 2021), whereas the latter only needs a few epochs (Zhang et al., 2022). Therefore, the speed of learning content modalities is significantly slower than learning categorical modalities. Joint training for multimodal recommendation is hard to deal with this imbalanced rate, introduces high computational overhead, and finally leads to an insufficient representation of items.
@@
Challenge 3: The influence of content modality to multimodal recommendation has not been fully studied. There are metrics designed for mLLMs to determine whether the aligned features are suitable for downstream tasks (Radford et al., 2021). However, the multimodal recommendation just incorporates vision and text features into recommendation and evaluates the final performance, which is indirect and does not consider their quality. If the vision or text features are not suitable, more effort is required for re-generation, decreasing the recommendation efficiency. We utilize a zero-shot evaluation in Sec. 4.4.2, and surprisingly find that the absence of vision features can gain a relatively 12.1% increase of recall@20. These observations call for us to directly evaluate the influence of multimodal features to recommendation and select suitable features.
@@
To manage the challenges above, we propose a multimodal recommendation framework AlignRec. For Challenge 1, we develop three alignment objectives for different contexts as illustrated in Fig. 1(d), respectively. The alignment within contents, namely inter-content alignment (ICA), aligns different content modality fields (e.g., vision and text) with the help of attention-based cross-modality encoder, and outputs a unified modality representation for one item. The alignment between content and categorical ID, namely content-category alignment (CCA), utilizes contrastive learning to bridge the gap between multimodal content features and user/item ID-based features for better utilization and adaptation of multimodal information. The user-item alignment (UIA) aligns users and their interacted items via cosine similarity. For Challenge 2, we propose to train AlignRec by first pre-training the inter-content alignment task, and then training the remaining two alignment tasks together with the recommendation goal. With the pre-training stage utilizing mLLM’s capacity to align content modality, the process of jointly learning multi-modality becomes more stable, which leads to a better item/user representation. For Challenge 3, we propose three intermediate evaluation protocols, including zero-shot, item-CF and mask modality recommendation, to directly evaluate whether the unified and aligned multimodal features generated via inter-content alignment are helpful for recommendation. It can help us select a better multimodal encoder and reduce the complexity of hyper-parameter search in training. The detailed differences on training and evaluation between AlignRec and current methods will be discussed in Fig. 2.
Our contributions can be concluded as follows:
@@
  • We also provide three classes of new intermediate protocols to evaluate whether the multimodal features are effective, help us select a better encoder, and accelerate the iteration of recommendation models. To the best of our knowledge, we are the first to intermediately evaluate the multimodal recommendation.
We conduct experiments on three public datasets and show the state-of-the-art performance of AlignRec compared to nine baselines. The intermediate evaluation with in-depth analysis also shows the superior performance of our alignments and training strategies. Our proceeded multimodal features show better performance than existing ones.
@@ 2. Related Work
Multimodal Recommendation. Multimodal recommendation is derived from collaborative filtering (CF) recommendation (Su and Khoshgoftaar, 2009), and incorporates multimodal information for better recommendation. It applies matrix factorization methods (Koren et al., 2009), BPR (Rendle et al., 2012) or LightGCN (He et al., 2020a) and is usually effective in cold-start or long-tail scenarios.
@@
Early methods concatenate vision or text features with ID-based features. VBPR (He and McAuley, 2016b) extends BPR-based recommendation by adding visual features. ACF (Chen et al., 2017) applies an attention layer to calculate modalities’ contributions to items. MMGCN (Wei et al., 2019) and GRCN (Wei et al., 2020) build modality graphs and use GCN (Kipf and Welling, 2016) to aggregate information. These methods ignore the misalignment between ID-based and content features, and thus cannot fully exploit multimodal information.
@@
Recently, some methods try to shorten the distance between item’s ID and content features. MMGCL (Yi et al., 2022) applies negative sampling to discover modalities’ relationship. SLMRec (Tao et al., 2022) builds self-supervised tasks to enhance the modality representation. BM3 (Zhou et al., 2023) builds text and vision alignment loss by contrastive view. They only focus on the content-category alignment and optimize all losses together, which is hard to converge with insufficient representations.
@@
There are also some methods trying to fuse vision and text information. LATTICE (Zhang et al., 2021) and FREEDOM (Zhou and Shen, 2023) construct modality-similarity graphs using multimodal features and combine these graphs with a pre-defined weight. MGCN (Yu et al., 2023) adopts attention layers to capture the modalities’ importance. These methods do not explicitly model the alignment, and will suffer from the modality incompatibility problem (e.g., conflict between text and vision).
@@
Vision-Language Pre-training. Recently, vision-language pre-training is effective in understanding vision and text information with large-scale data, which is helpful in multimodal recommendation. CLIP (Radford et al., 2021) proposes to align image and text via contrastive learning and pre-training. ALBEF (Li et al., 2021) designs an image-text contrastive learning task to align image and text features before fusing them. BLIP (Li et al., 2022) utilizes multi-task learning by building image-text pre-training tasks. BEiT3 (Wang et al., 2022) regards image as a foreign language and builds a general make-vision-language task to learn one integrated representation for multimodal features, with the help of multiway transformers (Bao et al., 2022). Due to the complicated structures of these models, many evaluation metrics are proposed. But they are missing in multimodal recommendation. Therefore, most of multimodal recommendation methods still use less powerful features.
Figure 2. An overview of AlignRec, where user ID, item ID, text and image are input, and user and item representations are output. A and B show the differences between AlignRec and current methods when training, where an intermediate evaluation module and a two-stage training strategy are proposed in AlignRec. C shows the overall architecture of AlignRec.
@@ 3. Problem Formulation
where, $u,i$ is a user-item pair satisfying $r_{u,i}=1$, $i^{\prime}$ is sampled from $\mathcal{I}$ satisfying $r_{u,i^{\prime}}=0$. $f:u\rightarrow\mathbb{R}^{d_{h}}$ and $g:i\rightarrow\mathbb{R}^{d_{h}}$ are mapping functions to map raw features (ID, images, text, etc.) into dense representations with dimension $d_{h}$; $s$ is a score function (e.g., inner product); $\sigma$ is the sigmoid function.
@@ 4.1. Framework Overview
The overview of AlignRec and its differences on training and evaluation from current methods are illustrated in Fig. 2. The whole framework consists of three alignments with an intermediate evaluation module. In this section, we will first introduce the architectures, then describe our three alignment objectives, and finally introduce training strategies and intermediate evaluation protocols.
@@ 4.2. Architecture Design
For generalization, AlignRec only contains three necessary modules: The multimodal encoder module $MMEnc$ aligns the vision and text modality knowledge into one unified representation. The aggregation module $Aggregator$ aggregates the neighbors’ information with $\mathcal{G}$. The fusion module $Fuser$ fuses multimodal and ID-based representations for recommendation. The goal is to produce item and user representations $\mathbf{h}_{i}$ and $\mathbf{h}_{u}$ for top-K recommendation.
@@ 4.2.1. Multimodal Encoder Module
It aims at aligning vision and text of the same item, and outputs a unified multimodal representation $\textbf{h}_{enc}^{i}$. We propose to adopt a transformer-based multimodal encoder to generate the unified multimodal representation. Formally, suppose $MMEnc$ is a multimodal encoder, and $h_{enc}^{i}$ is the unified representation for item $i$’s vision and text, then:
@@
In Eq. (2), the vision-text alignment is optimized through the cross-attention based transformers in $MMEnc$. We utilize the most advanced BEiT3 (Wang et al., 2022) as the backbone. The $MMEnc$ will align vision and text information via cross-modality experts (Bao et al., 2022). We train it in Sec. 4.3.1 through the mask-then-predict strategy for a better understanding of both vision and text modalities.
@@ 4.2.3. Fusion Module
The multimodal and ID-based hidden representations contain information of item/user from different fields or aspects. In AlignRec, we use a fusion module $Fuser$ to fuse this information into general item/user representations $\mathbf{h}^{i}$ and $\mathbf{h}^{u}$:
@@
Specifically, we use the element-wise addition operation in Eq. (8), and finally get a general representation for users and items. This module can also be replaced by more complicated fusion operations like attention in current methods.